OpenStudy (anonymous):

try factoring first, if to dificult use quadratice formula 1/x + 1/(x+5) = 1/3

OpenStudy (anonymous):

\[\frac{1}{x+5}+\frac{1}{x}=\frac{1}{3} \]\[\left\{x=\frac{1}{2} \left(1-\sqrt{61}\right),x=\frac{1}{2} \left(1+\sqrt{61}\right)\right\} \]\[\{-3.40512,4.40512\} \]

OpenStudy (anonymous):

not one of the answers?

OpenStudy (anonymous):

they are 11(+-)sqrt45 over 2 -1(+-)sqrt45 over 2 -11(+-)sqrt45 over 2 1(+-)sqrt45 over 2

OpenStudy (anonymous):

Move the RHS to the LHS and combine fractions:\[\frac{-x^2+x+15}{3 x (x+5)} \]\[\text{Solve}\left[-x^2+x+15==0,x\right]\to \left\{x\to \frac{1}{2} \left(1-\sqrt{61}\right),x\to \frac{1}{2} \left(1+\sqrt{61}\right)\right\} \]

OpenStudy (anonymous):

-11(+-)sqrt 45 over 2

OpenStudy (anonymous):

\[\frac{-11(+-) \sqrt{45} }{2} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{-11+ \sqrt{45} }{2}=-2.1459 ,\left(\frac{1}{x+5}+\frac{1}{x}\text{/.} x\to -2.1459\text{ }\right)=-0.115632 \]The answer should be equal to .333 or nearly so.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Thank you for the medal.